OpenStudy (anonymous):

–6 + 3x = –9 –1 –6 –5 –3

OpenStudy (anonymous):

you should post each question as a new post.

OpenStudy (anonymous):

1. 2x-26 = 10 bring the -26 to the right hand side 2x = 10 + 26 2x = 36 x = ......... you can do this. 2) -6 + 3x = -9 bring the -6 to the right hand side 3x = -3 x = ......... you can do this.

OpenStudy (kuoministers):

Nice teaching :)

OpenStudy (anonymous):

lol thanks.

OpenStudy (kuoministers):

hi @-Lexy-

OpenStudy (anonymous):

hello

OpenStudy (kuoministers):

2x = 36 x = .........

OpenStudy (anonymous):

18

OpenStudy (kuoministers):

there you go the answer is 18 gj

OpenStudy (anonymous):

yes that is correct, applying the opposite sign of operation.

OpenStudy (anonymous):

2 times x = 36 x = 36 divide by 2 x = 18

OpenStudy (anonymous):

i dont get number 2 though

OpenStudy (aonz):

post it as a diff question please

OpenStudy (anonymous):

2) -6 + 3x = -9 bring the -6 to the right hand side 3x = -3 3 times x = -3 x = -3 divide by 3 x = .........

OpenStudy (anonymous):

same principle is applied in question 2. and repost the rest of ur questions in new posts.

OpenStudy (anonymous):

1

OpenStudy (kuoministers):

no it is not 1 try again

OpenStudy (anonymous):

-3/3, simplify it down by cancelling the HCF/GCF

OpenStudy (anonymous):

it's not possible to be 9.

OpenStudy (anonymous):

u said 1 at first and that is partially correct but u are missing something.

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

nono, are u reading what I said.

OpenStudy (anonymous):

i dont get it

OpenStudy (kuoministers):

LOL WHATS -3 divided by 3

OpenStudy (kuoministers):

thats all it is :/

OpenStudy (anonymous):

lol.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

\[\frac{-3}{3} = \frac{-1}{1}=-1\]

OpenStudy (anonymous):

yes that is correct, because u can cancel the GCF/HCF of 3.

OpenStudy (kuoministers):

thats how easy it is...

OpenStudy (anonymous):

haha yeh, keep working on it @-Lexy-